OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (–2, 4); Slope: 5 A. y – 4 = 5(x + 2) B. y + 4 = 5(x + 2) C. y – 4 = 5(x – 2) D. y + 4 = 5x – 10

OpenStudy (anonymous):

Equation of line passing through one point is given by: \(y - y_1 = m(x - x_1)\) Here \(m\) is the slope and \((x_1, y_1) \) is your given point.. Put m = 5, \(x_1\) = -2 and \(y\) = 4 in above equation.. Go ahead..

OpenStudy (anonymous):

Keep \(x\) and \(y\) as such there..

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

Go through it once more..

OpenStudy (anonymous):

Remember, you are given x1 = -2 and not 2...

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

Put \(x_1\) here.. \(\large x - (x_1)\)

OpenStudy (anonymous):

Put value of \(x_1\) in that bracket..

OpenStudy (anonymous):

i still dont get it at all. i dont get how you find the answer. it doesnt make sense to me

OpenStudy (anonymous):

See you have one minus sign in front of \(x_1\) and \(x_1\) itself contains one minus sign ie \(\large x - (-2)\) So, \(-(-)\) becomes \(+\) It will become : \(y - 4 = 5(x + 2)\)

OpenStudy (anonymous):

what does x1 stand for? i dont get it?

OpenStudy (anonymous):

\((x_1, y_1)\) is the point from which the given line is passing.. See, you have given this point in your question ie \((-2, 4)\)..

OpenStudy (anonymous):

oooh ok. how do i know when to change the sign from minus to plus? like you gave me the equation (x - x1) but you said the answer is (x + 2) ?

OpenStudy (anonymous):

i mean you said now -(-) becomes + but i dont see how you came upon that it needed to be changed??

OpenStudy (anonymous):

See the equation is by default : (x - x1) Or you can write it as : (x - (x1)).. Right???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok i'm gonna try to do a problem on my own, is it ok if you tell me if I get the answer right?

OpenStudy (anonymous):

is it C?